oliviac1111:

What are the zeros of the quadratic function f(x) = 2x2 – 10x – 3?

xxemilianaxx:

Ok well first group the terms that have the same variable and move the constant to the other side of the problem

oliviac1111:

okay idk how to do that

xxemilianaxx:

\[f(x) = 2x2 - 10x - 3\]

xxemilianaxx:

Thats the problem now do you know how to do what i just said?

oliviac1111:

dont i have to combine the x's?

xxemilianaxx:

Take out the f(x) = part because we know it would equal 0

carmelle:

@xxemilianaxx wrote:
Ok well first group the terms that have the same variable and move the constant to the other side of the problem
try not to copy off of other website cus we aren't sure reliable they are. next time, explain the question in ur own words :)

xxemilianaxx:

so what would be the equation now?

xxemilianaxx:

@carmelle wrote:
@xxemilianaxx wrote:
Ok well first group the terms that have the same variable and move the constant to the other side of the problem
try not to copy off of other website cus we aren't sure reliable they are. next time, explain the question in ur own words :)
i havent even looked up the problem

oliviac1111:

can someone just give me the answer im freaking out here lol

oliviac1111:

or not argue about it

carmelle:

@xxemilianaxx wrote:
@carmelle wrote:
@xxemilianaxx wrote:
Ok well first group the terms that have the same variable and move the constant to the other side of the problem
try not to copy off of other website cus we aren't sure reliable they are. next time, explain the question in ur own words :)
i havent even looked up the problem
ok, I'll give u the benefit of the doubt, you may continue, I wanna see how u solve it

xxemilianaxx:

No because that's against the rules

@oliviac1111 wrote:
or not argue about it

oliviac1111:

u know what i mean

xxemilianaxx:

I can't just give you the answer's

xxemilianaxx:

That's why i'm going step by step wid you

oliviac1111:

ik i meant help

xxemilianaxx:

@xxemilianaxx wrote:
so what would be the equation now?
Answer this

oliviac1111:

ok it would be 2x^2-3

xxemilianaxx:

No, when you take away the f(x) = part it would result in just "2x2 – 10x – 3 = 0"

oliviac1111:

ohh

xxemilianaxx:

Now just take away the "=0" and rewrite the problem to equal 3

oliviac1111:

\[2\times2-10x=\]3

xxemilianaxx:

Perfect, Now can you factor the coefficient?

oliviac1111:

idk how to do that

xxemilianaxx:

Well do yk what x would equal

oliviac1111:

no..

oliviac1111:

got it

Florisalreadytaken:

can every comment before this be deleted please? thanks. so a quadratic funtion is in the form of \( 0=a^2+bx +c \) right? were given \( 0 = 2x^2 – 10x – 3 \) so we can determine that: \[ a=2 \] \ \[c-3 \] now, we can either do it by facoring, or by using the quadratic formula -- factoring is kinda eh, so imma do it using the quadratic formula: \[ \frac{-b\pm\sqrt{b^2-4ac}}{2a} \Rightarrow \frac{10\pm\sqrt{-10^2-4 \times (2 \times -3 )}}{2 \times 2} \] now that we have that, we got some math to do: \[\frac{10\pm\sqrt{-10^2-4 \times -6}}{4}\] \[ \frac{10\pm\sqrt{100+24}}{4}\] \[ \frac{10\pm\sqrt{124}}{4}\Rightarrow\frac{10\pm\sqrt{\cancel{124}}}{\cancel{4}} \] \[ \frac{10\pm\sqrt{62}}{2} \Rightarrow \frac{\cancel{10}\pm\sqrt{\cancel{62}}}{2} \] \[\frac{5\pm\sqrt{31}}{2} \] now there are 2 possible solutions -- we will once do it the negative way, and the 2nd x will be the positive one -- it doesnt matter which one is firs, and which one is second -- before i do that, do YOU know how to do that?

Florisalreadytaken:

@oliviac1111 ??

Florisalreadytaken:

\[\frac{5\pm\sqrt{31}}{2} \] \( \frac{5-\sqrt{31}}{2} \) AND \( \frac{5+\sqrt{31}}{2} \) \[ \frac{5-\sqrt{31}}{2} \] \[ \frac{5- 5.57 }{2} \] \[ \frac{-0.57 }{2} \] \[ x_1=-0.285 \] lets solve for the second root. \[ \frac{5+\sqrt{31}}{2} \] \[ \frac{5+ 5.57 }{2} \] \[ \frac{10.57 }{2} \] \[ x_2=5.285 \] and thats your answer.

oliviac1111:

thank u